OpenStudy (anonymous):

y=ax^2+c with a graph that includes (0,1/3) (-3, 82?3)

OpenStudy (anonymous):

(-3, 82/3)

OpenStudy (anonymous):

since \((0,\frac{1}{3})\) is on the graph you know \(c=\frac{1}{3}\) because that is what you get if \(x=0\)

OpenStudy (anonymous):

so you have \[y=ax^2+\frac{1}{3}\] put \(x=-3\) and \(y=\frac{82}{3}\) and solve for \(a\)

OpenStudy (anonymous):

in other words, solve \[\frac{82}{3}=9a+\frac{1}{3}\] for \(a\)

OpenStudy (anonymous):

3x^2+1/3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73 help..

OpenStudy (anonymous):

a=3

OpenStudy (nincompoop):

you can deduct 1/3 on both sides of the equation to isolate 9a show me what you came up with

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=-3x^2+2/3

OpenStudy (anonymous):

@nincompoop

OpenStudy (nincompoop):

this is what satellite left us: \[\frac{ 82 }{ 3 }=9a+\frac{ 1 }{3 }\]

OpenStudy (anonymous):

oh,

OpenStudy (nincompoop):

deducting 1/3 on both sides of the equation will leave us 81/3 = 9a now you just have to divide 9 on both sides of the equation so you are left with a

OpenStudy (anonymous):

y=3x^2-1/3

OpenStudy (anonymous):

@you. can you help meh?

OpenStudy (anonymous):

I'm 12 I don't know how to solve this

OpenStudy (anonymous):

oh

OpenStudy (nincompoop):

@selenamalter now just divide 81/3 by 9 to get the value of a \[\frac{ 81 }{ 3 }\div \frac{ 9 }{ 1 }\] we know that 9 x 9 is 81. so, 81/9 is 9 and 9/1 = 9 therefore, a = 9

OpenStudy (nincompoop):

sorry. THEREFORE, a = 9/3

OpenStudy (nincompoop):

9/3 = 3

OpenStudy (anonymous):

okay, so the full answer would be

OpenStudy (anonymous):

those are the answer options and D. none of these answers

OpenStudy (nincompoop):

assuming that where satellite left off is correct, then the answer would be D I trust satellite's math skills LOL

OpenStudy (anonymous):

;) wow, thanks :D

OpenStudy (nincompoop):

no problem gal

OpenStudy (anonymous):

i can go to bed now lol.... thank ya lad :D

OpenStudy (nincompoop):

give satellite the medal. I didn't really do anything. I just merely showed you what he wanted you to do.